jw21:

Car = V-8 sedan Years driven = first and second Miles driven = 15,500 miles first year and 14,500 miles second year The cost of depreciation will be $ .

Subshilava:

Notmeta is right 500 so 2 years and the miles driven so subtract the miles from each other 15,500-14,500/2= 1,000 1,000 divide the 2 years and you get 500.

RhianaB112:

The cost of depreciation will be (V-8) * 30,000(Vāˆ’8)āˆ—30,000 1Calculate the total miles driven over the two years: 15,500 + 14,500 = 30,000 miles 2Calculate the depreciation cost per mile: V-8 3Multiply the total miles driven by the depreciation cost per mile: (V-8) * 30,000(Vāˆ’8)āˆ—30,000

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

kikidoyoulovememwa:

REPLY 2 ME

toga:

To calculate the cost of depreciation, we can use the following formula: Cost of Depreciation = (Initial Value - Final Value) / Total Miles Driven Let's assume the initial value of the car is $30,000 and the final value after the second year is $20,000. Cost of Depreciation = ($30,000 - $20,000) / (15,500 + 14,500) = $10,000 / 30,000 = $0.33 per mile So, the cost of depreciation will be $0.33 per mile driven.